OpenStudy (kira_yamato):

When the function g(x) is divided by (x-3) and (x-2), the remainder is -10 and -8 respectively. Find the remainder when g(x) is divided by (x^2 - 5x + 6)

OpenStudy (anonymous):

See Question 2c of this exam. It's very similar.

OpenStudy (anonymous):

Let P(x)= (x-3)(x-2) Q(x) +ax+b

OpenStudy (anonymous):

P(3)=-10 ---> 3a+b =-10 P(2)= -8 --> 2a+b =-8

OpenStudy (anonymous):

First equation minus second equation a= -2 b=-8-2(-2) = -8+4 =-4

OpenStudy (anonymous):

So the answer is -2x-4